For Questions 14-16, use the following information. Suppose the annual expected mortality rate from coronary heart disease is 600 per 100,000 individuals.
14. In a specific community of 600,000 individuals, how many deaths are expected to be attributed to coronary heart disease over a one-year period?
15. The actual (observed) number of deaths attributed to coronary heart disease in the community was 4,500 individuals over a one-year period. What would be the observed/expected ratio for these deaths? Please reduce to the simplest form.
3
16. Of the 4,500 individuals that died from coronary heart disease, it was determined that twice the number of females died from it than males. How many males died from coronary heart disease in this community?
Answer #1

14. Expected deaths = (600000*600)/(100000) = 3600

15. Ratio = 4500 / 3600 = 5/4 = 1.25

16. Let no.of males died = x ; Then no.of females dies = 2x ; So x + 2x = 4500 ; Therefore x = 1500 ; So 1500 males died.

Please vote if this is useful.